A Complete Bibliography of Games: Research and Practice

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

10 May 2024 Version 1.00

Title word cross-reference

3 [12].

across [15]. Activities [29]. Advertising [33]. Affective [19]. AI [14]. Analysis [4]. Asia [27]. Audio [16, 12].

Back [10]. based [33]. Been [32]. Benefits [12]. Best [32]. Beyond [16, 13]. Box [25]. Broaden [22].

C [22]. Case [33]. Casual [15]. Cataclysmic [28]. Challenges [3, 29]. Charting [13]. Chat [21]. Citizens [14]. Co [31]. Co-Creating [31]. Cognition [15]. Combat [24]. Compassion [32]. Compliance [25]. Conscious [3]. Content [32, 12]. Control [24]. Counter [29]. Counter-extremism [29]. Creating [31]. creative [12]. Critical [15]. Cultures [18].

D [12]. Data [13]. Decentralized [17]. Deep [31]. Design [9, 19, 3, 31]. Developer [32]. Diegesis [20]. Digital [29]. Diversity [4]. Do [28].

East [27]. Editorial [1, 14]. Education [26]. Embracing [8]. Emotion [19]. Environmentally [3]. Epic [25]. Eudaimonic [20]. Exposure [24]. Extremism [23, 30, 29]. Extremist [29]. Extremity [4].

Faces [4]. First [14]. Framework [19]. Fusion [18]. Future [10, 17, 11, 13]. Futures [5, 31].

Game [28, 3, 16, 31, 13]. Gamer [18]. Games [2, 19, 5, 1, 33, 24, 26, 10, 17, 11, 14, 21, 15, 7, 6, 8, 4, 12]. Gaming [23, 12, 29, 27, 30]. Global [26]. Go [28, 17, 33].

Here [28]. Host [14].

Identity [18]. Illuminated [10]. Impact [10]. implementation [12]. Inaugural [1]. Individual [24]. Industry [23, 28]. Infrastructure [8]. Input [20]. Interaction [20]. Interactivity [19].

Labels [25]. Landscape [27]. Layoffs [28]. Lens [31]. Lessons [3]. Let [17]. Lifespan [15]. Lighthouse [1]. Location [33]. Location-based [33]. Locative [33]. Looking [10]. Loot [25].

Manifesto [2]. Many [4]. Massive [4]. Massive-scale [4]. Matter [6]. Matters [32]. May [14]. Meaning [20]. Meaningful [23]. Media [1]. Metaverse [11, 16, 21]. Microsoft [25]. Mobile [4]. Moderating [21]. Monetisation [4].

Needs [22]. Next [12]. Nexus [30]. Nintendo [25].

Observations [27]. Online [23, 11]. Opportunities [29].

pitfalls [12]. Platform [33]. Platforms [7, 25]. Play [18, 15]. Playable [1]. Player [20, 4]. point [12]. Pokémon [33]. Practice [9]. Practices [32]. Presence [25]. Probable [10]. Productive [2]. Psychological [2].

Reach [22]. Realism [12]. Reflection [20]. Reflections [27]. Research [2, 11]. Rightful [14]. Risks [18].

Safety [21]. scale [4]. Science [13]. Shopping [25]. Social [9, 10, 21, 7, 8]. Sociality [8]. Sony [25]. South [27]. Spaces [29]. Spending [4]. Standard [22]. Step [12]. Synthesis [15]. Systems [9].

Teaching [3]. technical [12]. Thriving [17]. Toxic [18]. Toxicity [24]. Transactional [4]. Transformative [31]. Transparency [23].

Understanding [4]. Unsatisfactory [25].

via [4]. Video [19, 16, 8]. view [12]. Voice [21].

Warned [32]. Warning [25]. Warnings [32]. Where [28]. World [11].

You’ve [32].

References

Deterding:2023:IEL

[1] Sebastian Deterding, Kenny Mitchell, Rachel Kowert, and Brad King. Inaugural editorial: a lighthouse for games and playable media. Games: Research and Practice, 1(1):1:1–1:??, March 2023. CODEN ???? ISSN 2832-

5516 (electronic). URL https://dl.acm.org/doi/10.1145/3585393.

Ballou:2023:MMP

[2] Nick Ballou. A manifesto for more productive psychological games research. Games: Research and Practice, 1(1):2:1–2:??, March 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3582929.

Fizek:2023:TEC

[3] Sonia Fizek, Mikhail Fiadotau, Hanna Wirman, and Maria Garda. Teaching environmentally conscious game design: Lessons and challenges. Games: Research and Practice, 1(1):3:1–3:??, March 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3583058.

Zendle:2023:MFM

[4] David Zendle, Catherine Flick, Sebastian Deterding, Joe Cutting, Elena Gordon-Petrovskaya, and Anders Drachen. The many faces of monetisation: Understanding the diversity and extremity of player spending in mobile games via massive-scale transactional analysis. Games: Research and Practice, 1(1):4:1–4:??, March 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3582927.

Deterding:2023:GF

[5] Sebastian Deterding, Kenny Mitchell, Rachel Kowert, and Brad King. Games futures I. Games: Research and Practice, 1(1):5:1–5:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.

org/doi/10.1145/3585394.

Taylor:2023:GM

[6] T. L. Taylor. Games matter. Games: Research and Practice, 1(1):6:1–6:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://

dl.acm.org/doi/10.1145/3582928.

Steinkuehler:2023:GSP

[7] Constance Steinkuehler. Games as social platforms. Games: Research and Practice, 1(1):7:1–7:??, March 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3582930.

Voll:2023:SVG

[8] Kimberly D. Voll. The sociality of video games: Embracing games as social infrastructure. Games: Research and Practice, 1(1):8:1–8:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/

doi/10.1145/3582931.

Cook:2023:DPS

[9] Daniel Cook. A design practice for social systems. Games: Research and Practice, 1(1):9:1–9:??, March 2023. CODEN ???? ISSN 2832-5516 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3583184.

Grace:2023:SIG

[10] Lindsay D. Grace. Social impact games, a probable future illuminated by looking back. Games: Research and Practice, 1(1):10:1–10:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/

doi/10.1145/3583986.

Koster:2023:OWM

[11] Raph Koster. From online world to metaverse: The future of online games and games research. Games: Research and Practice, 1(1):11:1–11:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.

org/doi/10.1145/3582932.

Rieger:2023:ANS

[12] Martin Rieger and Dominik Zingler. 3D audio: The next step in realism for gaming: Benefits and pitfalls of 3D audio in games from a creative content and technical implementation point of view. Games: Research and Practice, 1(1):12:1–12:??, March 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3583059.

Wallner:2023:BGC

[13] Günter Wallner and Anders Drachen. Beyond the game: Charting the future of game data science. Games: Research and Practice, 1(1):13:1– 13:??, March 2023. CODEN ???? ISSN 2832-5516 (electronic). URL

https://dl.acm.org/doi/10.1145/3582933.

Mitchell:2023:EGM

[14] Kenny Mitchell. Editorial: Games may host the first rightful AI citizens. Games: Research and Practice, 1(2):18:1–18:??, June 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3606834.

Squire:2023:CGC

[15] Kurt Squire, Garrison Wells, Maria J. Anderson-Coto, and Constance Steinkuehler. Casual games, cognition, and play across the lifespan: a critical synthesis. Games: Research and Practice, 1(2):14:1–14:??, June 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.

org/doi/10.1145/3594534.

McGregor:2023:VGA

[16] Iain McGregor. Video game audio in the metaverse and beyond. Games: Research and Practice, 1(2):15:1–15:??, June 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3593242.

Kajioka:2023:GLG

[17] E. Drake Kajioka. Games that let go: Thriving in a decentralized future. Games: Research and Practice, 1(2):16:1–16:??, June 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3600235.

Kowert:2023:YWY

[18] Rachel Kowert, Alexi Martel, and William B. Swann. You are what you play: The risks of identity fusion in toxic gamer cultures. Games: Research and Practice, 1(2):17:1–17:??, June 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3604402.

Croissant:2023:EDV

[19] Maximilian Croissant, Guy Schofield, and Cade McCall. Emotion design for video games: a framework for affective interactivity. Games: Research and Practice, 1(3):19:1–19:??, September 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3624537.

Gomez:2023:EPI

[20] Argenis Ramirez Gomez and Michael Lankes. Eudaimonic player interaction: a reflection on input diegesis and meaning. Games: Research and Practice, 1(3):20:1–20:??, September 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3617670.

Pappas:2023:SSG

[21] Mike Pappas. Social safety in games: Moderating voice chat in the metaverse. Games: Research and Practice, 1(3):21:1–21:??, September 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/

doi/10.1145/3624356.

Davidson:2023:CSN

[22] Guy Davidson. The C++ standard needs to broaden its reach. Games: Research and Practice, 1(3):22:1–22:??, September 2023. CODEN ????

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3626207.

Deedman:2023:EOM

[23] Jon Deedman. Extremism online: Meaningful transparency for the gaming industry. Games: Research and Practice, 1(4):23:1–23:??, December 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/

doi/10.1145/3630026.

Frommel:2023:ICE

[24] Julian Frommel and Regan L. Mandryk. Individual control over exposure to combat toxicity in games. Games: Research and Practice, 1(4):24:1– 24:??, December 2023. CODEN ???? ISSN 2832-5516 (electronic). URL

https://dl.acm.org/doi/10.1145/3633768.

Xiao:2023:SAL

[25] Leon Y. Xiao. Shopping around for loot box presence warning labels: Unsatisfactory compliance on Epic, Nintendo, Sony, and Microsoft platforms. Games: Research and Practice, 1(4):25:1–25:??, December 2023. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.

1145/3630631.

Grace:2023:GGE

[26] Lindsay D. Grace. Global games education now. Games: Research and Practice, 1(4):26:1–26:??, December 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3632364.

Tiwari:2023:ORG

[27] Siddhartha Paul Tiwari and Lindsay D. Grace. Observations and reflections on the gaming landscape in South East Asia. Games: Research and Practice, 1(4):27:1–27:??, December 2023. CODEN ???? ISSN 2832-5516

(electronic). URL https://dl.acm.org/doi/10.1145/3632286.

Farough:2024:GIC

[28] Amanda Farough and Michael Futter. The game industry’s cataclysmic layoffs: Where do we go from here? Games: Research and Practice, 2(1): 1:1–1:??, March 2024. CODEN ???? ISSN 2832-5516 (electronic). URL

https://dl.acm.org/doi/10.1145/3651280.

Schlegel:2024:EAD

[29] Linda Schlegel. Extremist activities in digital gaming spaces: Challenges and opportunities for counter-extremism. Games: Research and Practice, 2 (1):2:1–2:??, March 2024. CODEN ???? ISSN 2832-5516 (electronic). URL

https://dl.acm.org/doi/10.1145/3649810.

White:2024:NGE

[30] Jessica White and Galen Lamphere-Englund. At the nexus of gaming and extremism. Games: Research and Practice, 2(1):3:1–3:??, March 2024. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.org/

doi/10.1145/3643805.

Rusch:2024:CCT

[31] Doris C. Rusch. Co-creating transformative futures through the lens of deep game design. Games: Research and Practice, 2(1):4:1–4:??, March 2024. CODEN ???? ISSN 2832-5516 (electronic). URL https://dl.acm.

org/doi/10.1145/3643026.

Dunlap:2024:YBW

[32] Kelli N. Dunlap. You’ve (not) been warned: Developer best practices for content warnings and why compassion matters. Games: Research and Practice, 2(1):5:1–5:??, March 2024. CODEN ???? ISSN 2832-5516 (electronic).

URL https://dl.acm.org/doi/10.1145/3638287.

Dunham:2024:PGA

[33] John Dunham, Jiangnan Xu, Konstantinos Papangelis, Nicolas Lalone, Michael Saker, and David Schwartz. Pokémon GO as an advertising platform: The case for locative advertising in location-based games. Games: Research and Practice, 2(1):6:1–6:??, March 2024. CODEN ????

REFERENCES 8

ISSN 2832-5516 (electronic). URL https://dl.acm.org/doi/10.1145/
3641509.